rssLink RSS for all categories
 
icon_red
icon_green
icon_red
icon_red
icon_blue
icon_green
icon_green
icon_red
icon_red
icon_red
icon_orange
icon_green
icon_green
icon_green
icon_green
icon_blue
icon_green
icon_orange
icon_red
icon_green
icon_red
icon_red
icon_green
icon_red
icon_red
icon_red
icon_red
icon_orange
icon_green
 

FS#10380 — FS#14246 — TPIX

Attached to Project— Network
Incident
Backbone EU
In progress
0%
We detected some BGP flaps during the night and this morning at the TPIX.

We contacted them to have more information about the issue.

This task will be updated as soon as we know more.
Comment by OVH - Tuesday, 04 August 2015, 12:48PM

This is a strange issue. There was a loss of light at
about 1.30am CEST. During the night the traffic decreased
between 3.40am - 4.10am and 6.20-7.00am CEST. At about
9.30 we lost a BGP session (as hold-time
expired) and after this the traffic went back to normal.


Comment by OVH - Tuesday, 04 August 2015, 12:49PM

We've redirected the traffic to and from TPIX while we wait for IX to confirm that everything is fixed.


Comment by OVH - Wednesday, 05 August 2015, 15:41PM

We are redirecting the traffic to TPIX as IX confirmed that the issue has been resolved.